OpenStudy (anonymous):

What is the sum of the arithmetic sequence 140, 135, 130…, if there are 35 terms?

OpenStudy (anonymous):

do you know the formula

OpenStudy (anonymous):

isnt this one where you have to use both of the formulas ?

OpenStudy (anonymous):

that's better... then tell me both formulas :)

OpenStudy (anonymous):

an= a1 + (n-1)d sn= n/2(a1+an)

OpenStudy (anonymous):

like i know you have to use both formulas i just forget hich one you use first

OpenStudy (anonymous):

ah that's cool , so we just skipped a big part... now we need to know difference... do you know how to find it? if you have a1=140 a2=135 a3=130

OpenStudy (anonymous):

do yo use the first formula first ? the an= a1 + (n-1)d ?

OpenStudy (anonymous):

actually I think that we don't need to use first formula at all ! we're dealing only with \[\LARGE Sn=\frac n2[2a_1+(n-1)d]\] :) and we have a1 but we need to find "d"

OpenStudy (anonymous):

n= 35 ? right

OpenStudy (anonymous):

.. yes last term is a_35 and we only need to find difference , and we're done! \[\LARGE d=a_n-a_{n-1}\] \[\LARGE d=a_2-a_1\] ...=?

OpenStudy (anonymous):

i got 4934 for d.. i dont think thats right

OpenStudy (anonymous):

just use the formula I wrote above.. \[\LARGE d=a_2-a_1\] \[\LARGE d=135-140\] \[\LARGE d=?\]

OpenStudy (anonymous):

d is -5 ?

OpenStudy (anonymous):

yes it is... now substitute back into the formula: \[\LARGE Sn=\frac n2[2a_1+(n-1)d]\] \[\LARGE S_{35}=\frac {35}{2}[2\cdot 140+(35-1)\cdot (-5)]\]

OpenStudy (anonymous):

1925 ?

OpenStudy (anonymous):

That's correct, well done :)

OpenStudy (anonymous):

thank you again.. (:

OpenStudy (anonymous):

My pleasure :)